OpenStudy (anonymous):

Solve 2x^2 = 5x + 4 using the quadratic formula. Round your answer to the nearest tenth

OpenStudy (anonymous):

x1=-0.6 x2=3.1

OpenStudy (anonymous):

x1 = 1/4 [ 5 - sqrt(57) ] x2 = 1/4 [ 5 + sqrt(57) ]

OpenStudy (anonymous):

First you have to rearrange the equation to equal 0. 2x^2-5x-4=0 The from this form, you know that ax^2 + bx + c = 0 so a = 2, b = -5, c = -4 And then plug that into the quadratic formula \[X=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]